# Minimax rates in variance and covariance changepoint testing

### v1

@@ 6.1 Auxiliary Lemmas
The following Lemma can be viewed as a unification of Proposition 4.2 in Berthet and Rigollet (2013) and Theorem 1 in Koltchinskii and Lounici (2017).
@@
Fix any $p\in\mathbb{N}$ and $s\in[p]$, and let $X_{i}$ be centered and independent $p$-dimensional sub-Gaussian random variables with $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma$, for $i=1,\ldots,n$ and some $\Sigma\in\mathbb{R}^{p\times p}$. Assume further that $\left\lVert X_{i}\right\rVert_{\Psi_{2}}^{2}\leq u\left\lVert\Sigma\right\rVert_{\mathrm{op}}$ for all $i$ and some $u>0$. Let $\widehat{\Sigma}=n^{-1}\sum_{i=1}^{n}X_{i}X_{i}^{\top}$. There exists a constant $c_{0}>0$ depending only on $u$, such that, for all $x\geq 1$, we have
@@
The proof follows a similar strategy as the proof of Proposition 4.2 in Berthet and Rigollet (2013), which is a slightly less general result. We begin by claiming that there exist a subset $\mathcal{N}$ the unit sphere $S^{p-1}$ with cardinality at most (${p\choose s}9^{s}$ such that for any symmetric $A\in\mathbb{R}^{p\times p}$, we have
@@
where the last inequality follows from $A$ being symmetric, in which case operator norm agrees with the largest absolute eigenvalue of $A$. As in the proof of Theorem 4.4.5 in Vershynin (2018), as well as the results in Exercise 4.4.3 in the same book, there exists an $1/4$-net $\mathcal{N}_{I}$ of the unit sphere ${S}^{s-1}$ with cardinality at most $9^{s}$ such that
@@
where the first equality is due to Lemma 2.7.6 in Vershynin (2018). Since centering of a sub-Gamma random variable preserves the Orlicz-1 norm up to a universial constant, it follows that $\lVert v^{\top}(\widehat{\Sigma}-\Sigma)v\rVert_{\Psi_{1}}\lesssim u\left\lVert\Sigma\right\rVert_{\mathrm{op}}$. From Bernstein’s Inequality (Theorem 2.8.1 in Vershynin 2018), we thus have
@@
Let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. Assume further that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma$ for all $i$ and some positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. For any $s\in[p]$, let
$\widehat{\sigma}^{2}_{s}=\lambda_{\max}^{s}(\widehat{\Sigma}^{(1)})\wedge\lambda_{\max}^{s}(\widehat{\Sigma}^{(1)}),$ (232)
where $\widehat{\Sigma}^{(1)}=\gamma^{-1}\sum_{i=1}^{\gamma}X_{i}X_{i}^{\top}$, $\widehat{\Sigma}^{(2)}=\gamma^{-1}\sum_{i=n-\gamma+1}^{n}X_{i}X_{i}^{\top}$ and $\gamma=\gamma(p,n,s)=$
$\lceil s\log(ep/s)\vee\log\log(8n)\rceil$. For any $\delta\in(0,1)$, define the events
$\mathcal{E}_{5}$ $=\bigcap_{s\in\mathcal{S}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}(\Sigma)}\right\},$ (233)
@@
where $\mathcal{S}$ is given in (49), $r(p,n,s,t)$ is given in (43), $c_{4}=\delta^{2}(\delta^{2}+16\delta+64)^{-1}(2e\pi w^{2})^{-1}$, $c_{5}=4c_{0}\log(1+16/\delta)$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Then $\mathbb{P}\left(\mathcal{E}_{5}\cap\mathcal{E}_{6}\right)\geq 1-\delta/2$.
@@
$\mathbb{P}\left(\mathcal{E}_{5}^{\complement}\right)$ $=\mathbb{P}\left(\bigcup_{s\in\mathcal{S}}\left\{\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}(\Sigma)}\leq c_{4}\right\}\right)$ (235)
$\leq 2\sum_{s\in\mathcal{S}}\mathbb{P}\left(\lambda_{\max}^{s}\left(\gamma^{-1}\sum_{i=1}^{\gamma}X_{i}X_{i}^{\top}\right)\leq c_{4}\lambda_{\max}^{s}(\Sigma)\right)$ (236)
$\leq 2\sum_{s\in\mathcal{S}}\mathbb{P}\left(\sum_{i=1}^{\gamma}(v^{\top}X_{i})^{2}\leq\gamma c_{4}v^{\top}\Sigma v\right)$ (237)
$\leq 2\sum_{s\in\mathcal{S}}\exp\left[\frac{\gamma}{2}\left\{1+\log\left(2\pi w^{2}\right)-\log\left(\frac{1}{c_{4}}\right)\right\}\right]$ (238)
@@
Let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. For some $t\leq\lfloor n/2\rfloor$, assume that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{1}$ for $i\leq t$ , and $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{2}$ for $i\geq n-t+1$, where $\Sigma_{1}$ and $\Sigma_{2}$ are two positive definite matrices in $\mathbb{R}^{p\times p}$. Let $\widehat{\Sigma}_{1,t}=t^{-1}\sum_{i=1}^{t}X_{i}X_{i}^{\top}$ and $\widehat{\Sigma}_{2,t}=t^{-1}\sum_{i=n-t+1}^{n}X_{i}X_{i}^{\top}$, and for some $s\in\mathcal{S}$ define
$\widehat{\sigma}^{2}_{s}=\lambda_{\max}^{s}(\widehat{\Sigma}^{(1)})\wedge\lambda_{\max}^{s}(\widehat{\Sigma}^{(1)}),$ (253)
where $\gamma=\gamma(p,n,s)=\lceil s\log(ep/s)\vee\log\log(8n)\rceil$. For any $\delta\in(0,1)$, define the events
@@
$\mathbb{P}\left(\mathcal{E}_{7}^{\complement}\right)$$\leq\mathbb{P}\left\{\frac{\lambda_{\max}^{s}(\widehat{\Sigma}^{(1)})}{\lambda_{\max}^{s}(\Sigma_{1})}\geq c_{3}\right\}$ (256)
=$=\mathbb{P}\left\{\lambda_{\max}^{s}\left(\gamma^{-1}\sum_{i=1}^{\gamma}X_{i}X_{i}^{\top}\right)\geq c_{3}\lambda_{\max}^{s}(\Sigma_{1})\right\}$ (257)
$\leq\mathbb{P}\left\{\lambda_{\max}^{s}\left(\gamma^{-1}\sum_{i=1}^{\gamma}X_{i}X_{i}^{\top}-\Sigma\right)\geq(c_{3}-1)\lambda_{\max}^{s}(\Sigma_{1})\right\}.$ (258)
Now we apply Lemma 2 with $x=c_{0}^{-1}(c_{3}-1)\gamma$ and $c_{0}$ the constant from that Lemma to obtain
@@
Let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. Assume further that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma$ for all $i$ and some positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. For any $s\in[p]$, let
$\widehat{\sigma}^{2}_{s,\mathrm{con}}=\widehat{\lambda}^{s}_{\max}(\widehat{\Sigma}_{1,\lceil\beta\rceil})\wedge\widehat{\lambda}^{s}_{\max}(\widehat{\Sigma}_{2,\lceil\beta\rceil}),$ (265)
where $\widehat{\lambda}^{s}_{\max}(\cdot)$ is defined in (55), $\widehat{\Sigma}_{i,t}$ for $i=1,2$ is defined in (41) and $\beta=\beta(p,n,s)=s^{2}\{\log(ep)\vee\log\log(8n)\}$. For any $\delta\in(0,1)$, define the events
$\mathcal{E}_{9}$ $=\bigcap_{s\in\mathcal{S}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s,\mathrm{con}}}{\lambda_{\max}^{s}({\Sigma})}\right\},$ (266)
$\mathcal{E}_{10}$ $=\bigcap_{t\in\mathcal{T}}\bigcap_{s\in\mathcal{S}}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}})\leq c_{6}\lambda_{\max}^{s}({\Sigma})h(p,n,s,t)\right\},$ (267)
where $\mathcal{S}$ is given in (49), $h(p,n,s,t)$ is given in (60), $c_{4}=\delta^{2}(\delta^{2}+16\delta+64)^{-1}(2e\pi w^{2})^{-1}$, $c_{6}=c_{0}\{1+\log(4/\delta)\}$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Then $\mathbb{P}\left(\mathcal{E}_{9}\cap\mathcal{E}_{10}\right)\geq 1-\delta/2$.
@@
$\widehat{\lambda}_{\max}^{s}(A)\geq\lambda_{\max}^{s}(A)$ (268)
for any symmetric matrix $A$. It follows that
$\mathbb{P}(\mathcal{E}_{9})$ $=\mathbb{P}\left[\bigcap_{s\in\mathcal{S}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s,\mathrm{con}}}{\lambda_{\max}^{s}({\Sigma})}\right\}\right]$ (269)
$\geq\mathbb{P}\left[\bigcap_{s\in\mathcal{S}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}({\Sigma})}\right\}\right]$ (270)
$\geq 1-\delta/4,$ (271)
where $\widehat{\sigma}^{2}_{s}$ is as defined in Lemma 5, and the last inequality follows from the same arguments as in the proof of that Lemma.
@@
$\mathbb{P}(\mathcal{E}_{10}^{\complement})$ $\leq\sum_{t\in\mathcal{T}}\sum_{s\in\mathcal{S}}\mathbb{P}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}})>c_{6}\lambda_{\max}^{s}({\Sigma})h(p,n,s,t)\right\}$ (272)
@@
$\mathbb{P}(\mathcal{E}_{10}^{\complement})$ $\leq 2\sum_{t\in\mathcal{T}}\sum_{s\in\mathcal{S}}\sum_{j\in[p]}\sum_{k\in[p]}\mathbb{P}\left\{|Y_{1,t,j,k}|>\frac{c_{6}}{2s}\lambda_{\max}^{s}(\Sigma)h(p,n,s,t)\right\}.$ (276)
Due to Lemma 2.7.7 in Vershynin (2018) and Assumption 3, we have that
@@
$\leq u\lambda_{\max}^{1}(\Sigma)$ (278)
$\leq u\lambda_{\max}^{s}(\Sigma),$ (279)
@@
From Bernstein’s Inequality (Theorem 2.8.1 in Vershynin 2018), we thus have
$\mathbb{P}\left[|Y_{1,t,j,k}|\geq cu\lambda_{\max}^{s}(\Sigma)\left(\sqrt{\frac{x}{t}}\vee\frac{x}{t}\right)\right]$ $\leq 2e^{-x},$ (280)
@@
$\mathbb{P}\left\{|Y_{1,t,j,k}|>\frac{c_{6}}{2s}\lambda_{\max}^{s}(\Sigma)h(p,n,s,t)\right\}$ $\leq 2\exp\left[-\frac{8c_{6}}{c_{0}}\left\{\log(ep)\vee\log\log(8n)\right\}\right]$ (281)
@@
Let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. For some $t\leq\lfloor n/2\rfloor$, assume that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{1}$ for $i\leq t$ , and $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{2}$ for $i\geq n-t+1$, where $\Sigma_{1}$ and $\Sigma_{2}$ are two positive definite matrices in $\mathbb{R}^{p\times p}$ and $t\leq\lfloor n/2\rfloor$. For any fixed $s\in[p]$, let
$\widehat{\sigma}^{2}_{s,\mathrm{con}}=\widehat{\lambda}^{s}_{\max}(\widehat{\Sigma}_{1,\lceil\beta\rceil})\wedge\widehat{\lambda}^{s}_{\max}(\widehat{\Sigma}_{2,\lceil\beta\rceil}),$ (288)
where $\widehat{\lambda}^{s}_{\max}(\cdot)$ is defined in (55), $\widehat{\Sigma}_{i,t}$ for $i=1,2$ is defined in (41) and $\beta=\beta(p,n,s)=s^{2}\{\log(ep)\vee\log\log(8n)\}$. For any $\delta\in(0,1)$, define the events
$\mathcal{E}_{11}$ =$=\left\{\frac{\widehat{\sigma}^{2}_{s,\mathrm{con}}}{\lambda_{\max}^{s}(\Sigma_{1})\wedge\lambda_{\max}^{s}(\Sigma_{2})}\leq c_{7}\right\},$ (289)
$\mathcal{E}_{12}$ $=\bigcap_{i=1,2}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{i,t}-\Sigma_{i}})\leq\frac{c_{6}}{2}\lambda_{\max}^{s}({\Sigma_{i}})h(p,n,s,t)\right\},$ (290)
where $\mathcal{S}$ is given in (49), $h(p,n,s,t)$ is given in (60), $c_{6}=c_{0}\{1+\log(4/\delta)\}$, $c_{7}=1+2c_{0}\log(8/\delta)$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Then $\mathbb{P}\left(\mathcal{E}_{11}\cap\mathcal{E}_{12}\right)\geq 1-\delta/2$.

### v2

@@ 6 Auxiliary Lemmas
The following Lemma can be viewed as a unification of Proposition 4.2 in Berthet and Rigollet, 2013b and Theorem 1 in Koltchinskii and Lounici, (2017).
@@
Fix any $p\in\mathbb{N}$ and $s\in[p]$, and let $X_{i}$ be centered and independent $p$-dimensional sub-Gaussian random variables with $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma$, for $i=1,\ldots,n$ and some $\Sigma\in\mathbb{R}^{p\times p}$. Assume further that $\left\lVert X_{i}\right\rVert_{\Psi_{2}}^{2}\leq u\left\lVert\Sigma\right\rVert_{\mathrm{op}}$ for all $i$ and some $u>0$. Let $\widehat{\Sigma}=n^{-1}\sum_{i=1}^{n}X_{i}X_{i}^{\top}$ and let $\lambda_{\max}^{s}(\cdot)$ be defined as in (6). There exists a constant $c_{0}>0$ depending only on $u$, such that, for all $x\geq 1$, we have
@@
The proof follows a similar strategy as the proof of Proposition 4.2 in Berthet and Rigollet, 2013b , which is a slightly less general result. We begin by claiming that there exist a subset $\mathcal{N}$ the unit sphere $S^{p-1}$ with cardinality at most (${p\choose s}9^{s}$ such that for any symmetric $A\in\mathbb{R}^{p\times p}$, we have
@@
where the last inequality follows from $A$ being symmetric, in which case operator norm agrees with the largest absolute eigenvalue of $A$. As in the proof of Theorem 4.4.5 in Vershynin, (2018), as well as the results in Exercise 4.4.3 in the same book, there exists an $1/4$-net $\mathcal{N}_{I}$ of the unit sphere ${S}^{s-1}$ with cardinality at most $9^{s}$ such that
@@
where the first equality is due to Lemma 2.7.6 in Vershynin, (2018). Since centering of a sub-Gamma random variable preserves the Orlicz-1 norm up to a universial constant, it follows that $\lVert v^{\top}(\widehat{\Sigma}-\Sigma)v\rVert_{\Psi_{1}}\lesssim u\left\lVert\Sigma\right\rVert_{\mathrm{op}}$. From Bernstein’s Inequality (Theorem 2.8.1 in Vershynin, 2018), we thus have
@@
Let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. Assume further that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma$ for all $i\in[n]$ and some positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. Let $\widehat{\Sigma}_{1,t}=t^{-1}\sum_{i=1}^{t}X_{i}X_{i}^{\top}$ and $\widehat{\Sigma}_{2,t}=t^{-1}\sum_{i=n-t+1}^{n}X_{i}X_{i}^{\top}$. Define $\gamma(s)=\gamma(p,n,s)=s\log(ep/s)\vee\log\log(8n)$ and for any $s\in[p]$ such that $\gamma(s)\leq n$, let
$\widehat{\sigma}^{2}_{s}=\lambda_{\max}^{s}(\widehat{\Sigma}_{1,\lceil\gamma(s)\rceil})\wedge\lambda_{\max}^{s}(\widehat{\Sigma}_{2,\lceil\gamma(s)\rceil}),$ (235)
where $\lambda_{\max}^{s}(\cdot)$ is defined in (6). For any fixed $\delta\in(0,1)$, define the events
$\mathcal{E}_{5}$ $=\bigcap_{\begin{subarray}{c}s\in\mathcal{S};\\ \gamma(s)\leq n\end{subarray}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}(\Sigma)}\right\},$ (236)
@@
where $\mathcal{S}$ is given in (50), $\mathcal{T}$ is given in (21), $r(p,n,s,t)$ is given in (44), $c_{4}=\delta^{2}(\delta^{2}+16\delta+64)^{-1}(2e\pi w^{2})^{-1}$, $c_{5}=4c_{0}\log(1+16/\delta)$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Then $\mathbb{P}\left(\mathcal{E}_{5}\cap\mathcal{E}_{6}\right)\geq 1-\delta/2$.
@@
$\mathbb{P}\left(\mathcal{E}_{5}^{\complement}\right)$ $=\mathbb{P}\left(\bigcup_{\begin{subarray}{c}s\in\mathcal{S};\\ \gamma(s)\leq n\end{subarray}}\left\{\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}(\Sigma)}\leq c_{4}\right\}\right)$ (238)
$\leq 2\sum_{\begin{subarray}{c}s\in\mathcal{S};\\ \gamma(s)\leq n\end{subarray}}\mathbb{P}\left(\lambda_{\max}^{s}\left(\lceil\gamma(s)\rceil^{-1}\sum_{i=1}^{\lceil\gamma(s)\rceil}X_{i}X_{i}^{\top}\right)\leq c_{4}\lambda_{\max}^{s}(\Sigma)\right)$ (239)
$\leq 2\sum_{\begin{subarray}{c}s\in\mathcal{S};\\ \gamma(s)\leq n\end{subarray}}\mathbb{P}\left(\sum_{i=1}^{\lceil\gamma(s)\rceil}(v^{\top}X_{i})^{2}\leq\lceil\gamma(s)\rceil c_{4}v^{\top}\Sigma v\right)$ (240)
$\leq 2\sum_{s\in\mathcal{S}}\exp\left[\frac{\lceil\gamma(s)\rceil}{2}\left\{1+\log\left(2\pi w^{2}\right)-\log\left(\frac{1}{c_{4}}\right)\right\}\right]$ (241)
@@
Let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. For some $t\leq\lfloor n/2\rfloor$, assume that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{1}$ for $i\leq t$ , and $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{2}$ for $i\geq n-t+1$, where $\Sigma_{1}$ and $\Sigma_{2}$ are two positive definite matrices in $\mathbb{R}^{p\times p}$. Let $\widehat{\Sigma}_{1,t}=t^{-1}\sum_{i=1}^{t}X_{i}X_{i}^{\top}$ and $\widehat{\Sigma}_{2,t}=t^{-1}\sum_{i=n-t+1}^{n}X_{i}X_{i}^{\top}$, and let $\gamma(s)=\gamma(p,n,s)=s\log(ep/s)\vee\log\log(8n)$. For any fixed $s\in[p]$ such that $\gamma(s)\leq n$ define
$\widehat{\sigma}^{2}_{s}=\lambda_{\max}^{s}(\widehat{\Sigma}_{1,\lceil\gamma(s)\rceil})\wedge\lambda_{\max}^{s}(\widehat{\Sigma}_{2,\lceil\gamma(s)\rceil}),$ (256)
where $\lambda_{\max}^{s}(\cdot)$ is defined in (6). For the same $s$ and any $\delta\in(0,1)$, define the events
@@
$\mathbb{P}\left(\mathcal{E}_{7}^{\complement}\right)$$\leq\mathbb{P}\left\{\frac{\lambda_{\max}^{s}(\widehat{\Sigma}_{1,\lceil\gamma(s)\rceil})}{\lambda_{\max}^{s}(\Sigma_{1})}\geq c_{3}\right\}$ (259)
=$=\mathbb{P}\left\{\lambda_{\max}^{s}\left(\lceil\gamma(s)\rceil^{-1}\sum_{i=1}^{\lceil\gamma(s)\rceil}X_{i}X_{i}^{\top}\right)\geq c_{3}\lambda_{\max}^{s}(\Sigma_{1})\right\}$ (260)
$\leq\mathbb{P}\left\{\lambda_{\max}^{s}\left(\lceil\gamma(s)\rceil^{-1}\sum_{i=1}^{\lceil\gamma(s)\rceil}X_{i}X_{i}^{\top}-\Sigma\right)\geq(c_{3}-1)\lambda_{\max}^{s}(\Sigma_{1})\right\}.$ (261)
Now we apply Lemma 2 with $x=c_{0}^{-1}(c_{3}-1)\gamma(s)$ and $c_{0}$ the constant from that Lemma to obtain
@@
Assume $n\geq\log(ep)$ and let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. Assume further that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma$ for all $i\in[n]$ and some positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. Let $\widehat{\Sigma}_{1,t}=t^{-1}\sum_{i=1}^{t}X_{i}X_{i}^{\top}$ and $\widehat{\Sigma}_{2,t}=t^{-1}\sum_{i=n-t+1}^{n}X_{i}X_{i}^{\top}$ and let
$\widehat{\sigma}^{2}_{\mathrm{con}}$ $=\widehat{\lambda}^{1}_{\max}(\widehat{\Sigma}_{1,\lceil\log(ep)\rceil})\wedge\widehat{\lambda}^{1}_{\max}(\widehat{\Sigma}_{2,\lceil\log(ep)\rceil}),$ (268)
where $\widehat{\lambda}_{\max}^{s}(\cdot)$ is defined in (56). For any $\delta\in(0,1)$, define the events
$\mathcal{E}_{9}$ =$=\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{\mathrm{con}}}{\lambda_{\max}^{1}({\Sigma})}\right\},$ (269)
$\mathcal{E}_{10}$ $=\bigcap_{t\in\mathcal{T}}\bigcap_{s\in\mathcal{S}}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}})\leq c_{6}\lambda_{\max}^{1}({\Sigma})h(p,n,s,t)\right\},$ (270)
where $\mathcal{S}$ is given in (50), $\mathcal{T}$ is given in (21), $\lambda_{\max}^{s}(\cdot)$ is defined in (6), $h(p,n,s,t)$ is given in (62), $c_{4}=\delta^{2}(\delta^{2}+16\delta+64)^{-1}(2e\pi w^{2})^{-1}$, $c_{6}=c_{0}\{1+\log(4/\delta)\}$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Then $\mathbb{P}\left(\mathcal{E}_{9}\cap\mathcal{E}_{10}\right)\geq 1-\delta/2$.
@@
$\widehat{\lambda}_{\max}^{1}(A)\geq\lambda_{\max}^{1}(A)$ (271)
for any symmetric matrix $A$. By symmetry and the same arguments as in the proof of Lemma 5 it follows that
$\mathbb{P}(\mathcal{E}_{9})$ $=1-2\mathbb{P}\left\{{\lambda^{1}_{\max}(\widehat{\Sigma}_{1,\lceil\log(ep)\rceil})}<c_{4}\lambda_{\max}^{1}({\Sigma})\right\}$ (272)
$\geq 1-\delta/4.$ (273)
@@
$\mathbb{P}(\mathcal{E}_{10}^{\complement})$ $\leq\sum_{t\in\mathcal{T}}\sum_{s\in\mathcal{S}}\mathbb{P}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}})>c_{6}\lambda_{\max}^{1}({\Sigma})h(p,n,s,t)\right\}$ (274)
@@
$\mathbb{P}(\mathcal{E}_{10}^{\complement})$ $\leq 2\sum_{t\in\mathcal{T}}\sum_{s\in\mathcal{S}}\sum_{j\in[p]}\sum_{k\in[p]}\mathbb{P}\left\{|Y_{1,t,j,k}|>\frac{c_{6}}{2s}\lambda_{\max}^{1}(\Sigma)h(p,n,s,t)\right\}.$ (278)
Due to Lemma 2.7.7 in Vershynin, (2018) and Assumption 3, we have that
@@
$\leq u\lambda_{\max}^{1}(\Sigma),$ (280)
@@
From Bernstein’s Inequality (Theorem 2.8.1 in Vershynin, 2018), we thus have
$\mathbb{P}\left[|Y_{1,t,j,k}|\geq cu\lambda_{\max}^{1}(\Sigma)\left(\sqrt{\frac{x}{t}}\vee\frac{x}{t}\right)\right]$ $\leq 2e^{-x},$ (281)
@@
$\mathbb{P}\left\{|Y_{1,t,j,k}|>\frac{c_{6}}{2s}\lambda_{\max}^{1}(\Sigma)h(p,n,s,t)\right\}$ $\leq 2\exp\left[-\frac{8c_{6}}{c_{0}}\left\{\log(ep)\vee\log\log(8n)\right\}\right]$ (282)
@@
Assume $n\geq\log(ep)$ and let $X_{1},\ldots,X_{n}$ be $p$-dimensional random variables satisfying Assumption 3 for some $w,u>0$. For some $t\leq\lfloor n/2\rfloor$, assume that $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{1}$ for $i\leq t$ , and $\mathbb{E}X_{i}X_{i}^{\top}=\Sigma_{2}$ for $i\geq n-t+1$, where $\Sigma_{1}$ and $\Sigma_{2}$ are two positive definite matrices in $\mathbb{R}^{p\times p}$. Let $\widehat{\Sigma}_{1,t}=t^{-1}\sum_{i=1}^{t}X_{i}X_{i}^{\top}$ and $\widehat{\Sigma}_{2,t}=t^{-1}\sum_{i=n-t+1}^{n}X_{i}X_{i}^{\top}$ and let
$\widehat{\sigma}^{2}_{\mathrm{con}}$ $=\widehat{\lambda}^{1}_{\max}(\widehat{\Sigma}_{1,\lceil\log(ep)\rceil})\wedge\widehat{\lambda}^{1}_{\max}(\widehat{\Sigma}_{2,\lceil\log(ep)\rceil}),$ (289)
where $\widehat{\lambda}_{\max}^{s}(\cdot)$ is defined in (56). For any fixed $\delta\in(0,1)$ and $s\in[p]$, define the events
$\mathcal{E}_{11}$ =$=\left\{\frac{\widehat{\sigma}^{2}_{\mathrm{con}}}{\lambda_{\max}^{1}(\Sigma_{1})\wedge\lambda_{\max}^{1}(\Sigma_{2})}\leq c_{3}\right\},$ (290)
$\mathcal{E}_{12}$ $=\bigcap_{i=1,2}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{i,t}-\Sigma_{i}})\leq\frac{c_{6}}{2}\lambda_{\max}^{1}({\Sigma_{i}})h(p,n,s,t)\right\},$ (291)
where $h(p,n,s,t)$ is given in (62), $\lambda_{\max}^{s}(\cdot)$ is defined in (6), $c_{3}=1+c_{0}\log(1+4/\delta)$, $c_{6}=c_{0}\{1+\log(4/\delta)\}$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Then $\mathbb{P}\left(\mathcal{E}_{11}\cap\mathcal{E}_{12}\right)\geq 1-\delta/2$.